# CopyJudge: Automated Copyright Infringement Identification and Mitigation in Text-to-Image Diffusion Models

### v1

@@
Abstract
Assessing whether AI-generated images are substantially similar to copyrighted works is a crucial step in resolving copyright disputes. In this paper, we propose CopyJudge, an automated copyright infringement identification framework that leverages large vision-language models (LVLMs) to simulate practical court processes for determining substantial similarity between copyrighted images and those generated by text-to-image diffusion models. Specifically, we employ an abstraction-filtration-comparison test framework with multi-LVLM debate to assess the likelihood of infringement and provide detailed judgment rationales. Based on the judgments, we further introduce a general LVLM-based mitigation strategy that automatically optimizes infringing prompts by avoiding sensitive expressions while preserving the non-infringing content. Besides, our approach can be enhanced by exploring non-infringing noise vectors within the diffusion latent space via reinforcement learning, even without modifying the original prompts. Experimental results show that our identification method achieves comparable state-of-the-art performance, while offering superior generalization and interpretability across various forms of infringement, and that our mitigation method could more effectively mitigate memorization and IP infringement without losing non-infringing expressions.
Machine Learning, ICML
1 Introduction
Text-to-image generative models (Rombach et al., 2022; Betker et al., 2023; Team et al., 2023; Esser et al., 2024; Hurst et al., 2024; Zhang et al., 2023b, a; Hintersdorf et al., 2024) have transformed creative industries by producing detailed visuals from text prompts. However, these models have been found to sometimes memorize and reproduce content from their training data (Carlini et al., 2023; Somepalli et al., 2023a; Ren et al., 2024; Wang et al., 2024c; Shi et al., 2024b, a; Zhang et al., 2024). This raises significant concerns about copyright infringement, especially when the generated images closely resemble existing copyrighted works. According to U.S. law (rot, 1970), also referenced by most countries, a work can be considered infringing if it constitutes substantial similarity to another work  [^1] Determining infringement also requires proving that the infringer had access to the copyrighted content, which is not the focus of our discussion in this paper.. Therefore, determining whether AI-generated images infringe on copyright requires a clear and reliable method to compare them with copyrighted materials to identify substantial similarity.
@@ 1 Introduction
Recently, large-scale models have already been successfully applied as judges in fields such as finance, education, and healthcare (Gu et al., 2024; Li et al., 2024; Zhuge et al., 2024). In this paper, we attempt to leverage large vision-language models (LVLMs) to model the practical court decisions on substantial similarity. However, directly applying large models for infringement identification may face unreliable outputs due to their limited comprehension or potential misinterpretation. To address this, we propose CopyJudge, an automated abstraction-filtration-comparison framework with multi-LVLM debate to reliably follow the court decision process on identifying substantial similarity.
Specifically, referring to the software abstraction test (Abramson, 2002), we decompose the image into different layers or elements, such as composition, color, and theme, to distinguish between the basic concepts of the image and its specific expressions. Then, we filter out parts that are not copyright-protected, such as public and functional expressions. Finally, we compare the filtered portions to assess whether there is substantial similarity. To enhance the reliability of the judgment, we employ a multi-agent debate (Du et al., 2023; Chan et al., 2023) method where multiple LVLMs discuss and score the similarity. Each LVLM can make judgments based on the scores and reasons provided by other LVLMs. Ultimately, another LVLM-based meta-judge gives the final score and rationale based on the consensus of the debate. To enhance the consistency with human preferences, we inject human priors into each LVLM via few-shot demonstrations (Agarwal et al., 2024).
Given the judging results, we further explore how to mitigate infringement issues in text-to-image diffusion models. Utilizing our CopyJudge, we propose an automated black-box infringement mitigation strategy that leverages a defense LVLM to iteratively optimize the input infringing prompts. This process avoids generating sensitive infringing expressions by querying supporting infringement rationales, while preserving the integrity of the original content. Moreover, if the input latent of the diffusion model is controllable, we could further enhance the mitigation approach by exploring specific non-infringing noise vectors within the latent space in a reinforcement manner, with the reward being reducing the predicted infringement score. This helps avoid infringement while maintaining the desired output characteristics, even without changing the original prompts.
@@ 2 Related Work
Copyright infringement in text-to-image models. Recent research (Carlini et al., 2023; Somepalli et al., 2023a, b; Gu et al., 2023; Wang et al., 2024c; Wen et al., 2024; Chiba-Okabe, 2024) highlights the potential for copyright infringement in text-to-image models. These models are trained on vast datasets that often include copyrighted material, which could inadvertently be memorized by the model during training (Vyas et al., 2023; Ren et al., 2024). Additionally, several studies have pointed out that synthetic images generated by these models might violate IP rights due to the inclusion of elements or styles that resemble copyrighted works (Poland, 2023; Wang et al., 2024a). Specifically, models like stable diffusion (Rombach et al., 2022) may generate images that bear close resemblances to copyrighted artworks, thus raising concerns about IP infringement (Shi et al., 2024b).
@@
Image infringement detection and mitigation. The current mainstream infringing image detection methods primarily measure the distance or invariance in pixel or embedding space (Carlini et al., 2023; Somepalli et al., 2023a; Shi et al., 2024b; Wang et al., 2021, 2024b). For example, Carlini et al. uses the $L_{2}$ norm to retrieve memorized images. Somepalli et al. use SSCD (Pizzi et al., 2022), which learns the invariance of image transformations to identify memorized prompts by comparing the generated images with the original training ones. Zhang et al. compare image similarity using the LPIPS distance, which aligns with human perception but has limitations in capturing certain nuances. Wang et al. transform the replication level of each image replica pair into a probability density function. Studies (Wen et al., 2024; Wang et al., 2024d) have shown that these methods have lower generalization capabilities and lack interpretability because they do not fully align with human judgment standards. For copyright infringement mitigation, the current approaches mainly involve machine unlearning to remove the model’s memory of copyright information (Bourtoule et al., 2021; Nguyen et al., 2022; Kumari et al., 2023; Zhang et al., 2024) or deleting duplicated samples from the training data (Webster et al., 2023; Somepalli et al., 2023b). These methods often require additional model training. On the other hand, Wen et al. have revealed the overfitting of memorized samples to specific input texts and attempt to modify prompts to mitigate the generation of memorized data. Similarly, Wang et al. leverage LVLMs to detect copyright information and use this information as negative prompts to suppress the generation of infringing images.
3 LVLM for Infringement Identification
@@
Figure 1: CopyJudge: an automated abstraction-filtering-comparison framework for image copyright infringement identification. This LVLM-based framework automatically decides whether a generated image infringes copyright by combining copyright expression extraction with infringement determination through multi-LVLM debate.
3.1 Problem Formulation
Our goal is to determine whether an image infringes the copyright of a known copyrighted image. Based on U.S. law (rot, 1970) and similar laws in other countries, given an image $x$ created with access to the copyrighted image $x_{cr}$, if $x$ and $x_{cr}$ exhibit substantial similarity, then $x$ is deemed to infringe the copyright of $x_{cr}$.
@@ 3.1 Problem Formulation
Motivated by this, we aim to establish a substantial similarity identification model $f$, which takes $x$ and $x_{cr}$ as inputs and outputs a similarity score $s$. When $s$ exceeds a threshold $\gamma$, we determine that $x$ infringes on $x_{cr}$. This can be defined as:
$\text{IsInfringement}(x)=\mathbb{I}(f(x,x_{{cr}})>\gamma),$ (1)
@@
3.2 Abstraction-Filtering-Comparison Framework
@@
$\pi_{i}(x,x_{{cr}},z^{c},z^{c}_{cr},p_{i},D_{h},S_{h})\to(s_{i},c_{i},r_{i}).$ (4)
$\left|s_{i}-s_{j}\right|\leq\alpha\quad\forall i,j\in\{1,2,\dots,N\}.$ (5)
$\pi_{f}(x,x_{{cr}},z^{c},z^{c}_{cr},p_{f},D_{h},S_{h},S_{m},C_{m},R_{m})\to(s_{f},c_{f},r_{f}),$ (6)
where $S_{m}$, $C_{m}$, and $R_{m}$ represent the set of scores, confidence levels, and rationales from $N$ LVLMs after reaching consensus in the $m$-th ($m\leq M$) debate. By combining the strengths of individual agents and iterative debating, the approach could achieve a reliable assessment of visual similarity. Furthermore, we can determine whether the generated image constitutes infringement based on whether the final similarity score exceeds a specific threshold $\gamma$:
$\text{IsInfringement}(x)=\mathbb{I}(s_{f})>\gamma.$ (7)
The whole two-stage process ensures a comprehensive and reliable evaluation by integrating multiple perspectives and rigorous analysis. The complete algorithm and instruction prompts can be found in appendix A.1 and A.2.
4 LVLM for Infringement Mitigation
Figure 2: Copyright infringement mitigation framework with controlling input prompts and latent noise. For prompt control, the defense agent iteratively optimizes prompts to mitigate copyright risks in generated images given the CopyJudge’s feedback. For latent control, it integrates a RL-based defense agent with reward-guided latent sampling to reduce the predicted infringement scores.
4.1 Mitigation via LVLM-based Prompt Control
$\pi_{p}(x^{t},x_{{cr}},p^{t},p^{c},s^{t}_{f},c^{t}_{f},r^{t}_{f})\to p^{t+1}.$ (8)
Here, $p^{c}$ requires the modifier to alter the infringing expression while preserving the original expression as much as possible to avoid generating meaningless images. This mitigation strategy does not require any knowledge of text-to-image models, making it suitable for general black-box scenarios. The algorithm and instruction prompts can be found in appendix B.1 and B.2.
4.2 Mitigation via RL-based Latent Control
$R(z)=-\log(s_{f}).$ (9)
$L(\omega)=\mathbb{E}_{z\sim\pi_{\omega}}[R(z)].$ (10)
$\nabla_{\omega}L(\omega)=\mathbb{E}_{z\sim\pi_{\omega}}[\nabla_{\omega}\log(\pi_{\omega})R(z)].$ (11)
$z^{\prime}=z+\beta\epsilon,\quad\epsilon\sim\pi_{\omega}(z),$ (12)
where $\beta$ is the step size. We further conduct normalization for the latent variables to maintain stability and prevent extreme deviations. This RL-based approach allows the agent to explore variations in the latent space, thereby improving its ability to generate non-infringing content. The detailed algorithm can be found in appendix B.1.
5 Experiments
5.1 Infringement Identification Experiments
Dataset. Firstly, we use D-Rep dataset (Wang et al., 2024b), which contains realistic and generated image pairs scored by human from 0 to 5 according to their similarity. Following its setting, we consider samples with scores of 4 or above as infringement samples, while the rest were non-infringement samples. We use the 4,000 official test images. In addition, we also consider the specific IP infringement. Referring to (Ma et al., 2024), we select 10 well-known cartoon characters and 10 artworks from Wikipedia (all can be found in B.2). 3 different text-to-image models—Stable Diffusion v2, Kandinsky2-2 (Razzhigaev et al., 2023), and Stable Diffusion XL (Podell et al., 2023)—are used to generate images. For each item, we manually select one infringing image and one non-infringing image from each model, resulting in a total of 60 positive samples and 60 negative samples.
Baselines. We select 4 commonly used distance-based image copy detection metrics: $L_{2}$ norm (Carlini et al., 2023), which directly measures pixel-wise differences; LPIPS (Zhang et al., 2018), which captures perceptual similarity based on deep network features; SSCD (Pizzi et al., 2022), which learns transformation-invariant representations to match generated images with their original training counterparts; and RLCP (Shi et al., 2024b), which integrates semantic and perceptual scores for a more comprehensive similarity assessment. Additionally, we compare our approach with the state-of-the-art image copy detection method PDF-Emb (Wang et al., 2024b), which models the similariy level of image pairs as a probability density function. We implement all methods using their official open-source code.
Implementation detail. Although our approach does not rely on a specific LVLM, we choose GPT-4o (Hurst et al., 2024) as our agent by default. As for multi-agent debate, we use 3 agents with a maximum of 5 iterations. We use random 3 images from each level (0-5) in the D-Rep training set as human priors to present to the agent.
Results. From the results in Table 1, it is evident that traditional image copy detection methods exhibit limitations in the copyright infringement identification task. Our approach significantly outperforms most methods. For the state-of-the-art method, PDF-Emb, which was trained on 36,000 samples from the D-Rep, our performance on D-Rep is slightly inferior. However, its poor performance on the Cartoon IP and Artwork dataset highlights its lack of generalization capability, whereas our method demonstrates equally excellent results across datasets. Figure LABEL:fig:res1 illustrates the prediction score distributions for all methods, showing that our approach achieves a relatively more distinct boundary between infringing and non-infringing cases. More results and details on time cost can be found in appendix A.3 and A.4.
[…]
Method / Dataset D-Rep Cartoon Artwork
$L_{2}$ norm 27.42 / 42.32 51.67 / 67.42 50.00 / 66.67
LPIPS 26.85 / 42.29 56.67 / 68.29 66.67 / 67.74
SSCD 43.65 / 44.37 58.33 / 70.59 90.00 / 90.32
RLCP 27.02 / 42.23 50.00 / 66.67 50.00 / 66.67
PDF-Emb 79.90 / 57.10 61.67 / 71.60 81.67 / 80.70
CopyJudge (Ours) 75.67 / 51.04 91.67 / 90.91 93.33 / 92.86
5.2 Infringement Mitigation Experiments
5.2.1 Memorization Mitigation
Figure 5: Visualization of generated images and corresponding prompts before and after memorization mitigation.
Table 2: Memorization mitigation results.
Method Infringement Score $\downarrow$ CLIP Score $\uparrow$
Raw 0.869 31.65
Wen et al.’s Method 0.605 28.96
PC (ours) 0.468 28.31
Figure 6: Visualization of generated images before and after IP infringement mitigation.
5.2.2 IP Infringement Mitigation
For explicit infringement, we validate both prompt control (PC) and latent control (LC). For implicit infringement, due to the commercial model DALL·E’s inability to customize latents, we only evaluate prompt control. In both scenarios, we compare our approach with the only prompt-based IP infringement mitigation method (Wang et al., 2024d), which detects potentially infringed works using LVLM and inputs the detected work information as a negative prompt (Neg-P) into the generative model to avoid infringement. We follow the same settings as in the original paper, but the paper does not address how to handle commercial models that cannot accept negative prompts. For such models, we simply appended the suffix ”without [copyrighted work information (e.g., name, author, etc.)]” to the original prompt to simulate a negative prompt.
Table 3: Explicit IP infringement mitigation results on Stable Diffusion v2 (Infringement Score $\downarrow$ / CLIP Score $\uparrow$).
Method / Dataset Cartoon Artwork
Raw 0.810 / 32.39 0.760 / 35.21
Neg-P 0.597 / 30.91 0.616 / 32.07
LC (Ours) 0.673 / 33.19 0.665 / 33.50
PC (Ours) 0.450 / 29.48 0.417 / 29.46
PC + LC (Ours) 0.350 / 28.67 0.353 / 28.82
Table 4: Implicit IP infringement mitigation results on DALL·E 3 (Infringement Score $\downarrow$ / CLIP Score $\uparrow$).
Method / Dataset Cartoon Artwork
Raw 0.783 / 30.54 0.697 / 29.80
Neg-P 0.710 / 29.27 0.689 / 29.34
PC (Ours) 0.547 / 27.92 0.431 / 28.81
Results. From Tables 3 and 4, it can be seen that our method significantly reduces the likelihood of infringement, both for explicit and implicit infringement, with only a slight drop in CLIP Score. The infringement score after only latent control is relatively higher than after prompt control because retrieving non-infringing latents without changing the prompt is quite challenging. However, we can still effectively reduce the infringement score while maintaining higher image-text matching quality. Figure 6 shows visualization results, where it can be observed that we avoid the IP infringement while preserving user requirements. Additional results and detailed time cost analysis are provided in Appendix B.3 and B.4.
Table 5: Ablation studies on different identification modules.
LVLM AFC MAD DEM ACC / F1-Score $\uparrow$
- - - 73.70 $\pm$ 0.68 / 42.44 $\pm$ 1.08
- - 76.50 $\pm$ 1.70 / 44.79 $\pm$ 2.30
- - 75.00 $\pm$ 1.14 / 43.48 $\pm$ 1.57
- - 76.50 $\pm$ 1.14 / 48.35 $\pm$ 1.53
- 77.50 $\pm$ 1.70 / 47.62 $\pm$ 1.81
78.50 $\pm$ 2.00 / 51.70 $\pm$ 2.19
5.3 Ablation Studies
6 Limitations and Future Work
While our method shows promise, its performance is currently constrained by the availability of labeled data, making it challenging to fully evaluate its alignment with human judgment. To address this, we believe that developing a more comprehensive dataset containing detailed human judgment criteria is crucial. In addition, we are looking forward to stronger attacks to test the robustness of our mitigation strategies. Improving the retrieval of infringing images will also provide valuable insights into the relationship between generative models and copyrighted content. Lastly, although our approach does not require a specific LVLM, we plan to explore other models beyond GPT in advancing copyright protection in the future.
7 Conclusion
A.1 Algorithm
In Algorithm 1, we present the detailed procedure for automated infringement identification based on the abstraction-filtering-comparison framework. The algorithm systematically extracts copyrightable expressions from images, filters out non-protectable elements, and employs a multi-agent debate among LVLMs to assess substantial similarity.
@@ A.1 Algorithm
23:  $\text{IsInfringement}(x)\leftarrow\mathbb{I}(s_{f}>\gamma)$
A.2 Instructive Prompts Used
    Agent 1 score: S1, confidence: C1, reason: R1
    Agent 2 score: S2, confidence: C2, reason: R2
    ……(if any)
    Scores: [S1, S2, S3, ……(if any)]
    Confidences: [C1, C2, C3, ……(if any)]
    Rationales: [R1, R2, R3, ……(if any)]
A.3 Additional Results
Figure 7: Successful copyright infringement identification by CopyJudge.
Figure 8: Failure case of copyright infringement identification by CopyJudge.
A.4 Time Consumption
The computational cost of running the CopyJudge is an important consideration. Table 6 provides the average runtime per sample for different steps in the process. We call the official API for testing and find that a single API call, along with extracting useful information, takes approximately 12s. The main time overhead comes from the multi-agent debate process. We acknowledge that this could be a limitation of our approach, while it allows for higher accuracy. In the future, we aim to improve the deployment and interaction of LVLMs to achieve faster decision-making.
Process Step Average Time (seconds)
Single GPT-4o 12.3
Copyright Expression Extraction (Abstraction + Filtration) 13.1
Whole Process 88.7
Table 6: Time analysis of the infringement identification process.
Appendix B Supplementary Information on Infringement Mitigation
B.1 Algorithm
For infringement mitigation, we use two strategies: prompt control and latent control, corresponding to Algorithm 2 and Algorithm 3, respectively.
B.2 Instructive Prompts and IP Used
    Original prompt: [Original Prompt]
    Original prompt: [Original Prompt]

### v2

@@
Abstract.
Assessing whether AI-generated images are substantially similar to copyrighted works is a crucial step in resolving copyright disputes. In this paper, we propose CopyJudge, an automated copyright infringement identification framework that leverages large vision-language models (LVLMs) to simulate practical court processes for determining substantial similarity between copyrighted images and those generated by text-to-image diffusion models. Specifically, we employ an abstraction-filtration-comparison test framework with multi-LVLM debate to assess the likelihood of infringement and provide detailed judgment rationales. Based on the judgments, we further introduce a general LVLM-based mitigation strategy that automatically optimizes infringing prompts by avoiding sensitive expressions while preserving the non-infringing content. Besides, our approach can be enhanced by exploring non-infringing noise vectors within the diffusion latent space via reinforcement learning, even without modifying the original prompts. Experimental results show that our identification method achieves comparable state-of-the-art performance, while offering superior generalization and interpretability across various forms of infringement, and that our mitigation method could more effectively mitigate memorization and IP infringement without losing non-infringing expressions [^1] Code is available at https://github.com/shunchang-liu/CopyJudge..
Copyright Protection, Text-to-image Diffusion Model
copyright: nonejournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755027isbn: 979-8-4007-2035-2/2025/10ccs: Security and privacy Social aspects of security and privacyccs: Computing methodologies Machine learning
1. Introduction
Text-to-image generative models (Rombach et al., 2022; Betker et al., 2023; Team et al., 2023; Esser et al., 2024; Hurst et al., 2024; Zhang et al., 2023a, b; Hintersdorf et al., 2024) have transformed creative industries by producing detailed visuals from text prompts. However, these models have been found to sometimes memorize and reproduce content from their training data (Carlini et al., 2023; Somepalli et al., 2023a; Ren et al., 2024; Wang et al., 2024d; Shi et al., 2024b, a; Zhang et al., 2024). This raises significant concerns about copyright infringement, especially when the generated images closely resemble existing copyrighted works. According to U.S. law (rot, 1970), also referenced by most countries, a work can be considered infringing if it constitutes substantial similarity to another work. Therefore, determining whether AI-generated images infringe on copyright requires a clear and reliable method to compare them with copyrighted materials to identify substantial similarity.
@@ 1. Introduction
Recently, large-scale models have already been successfully applied as judges in fields such as finance, education, and healthcare (Gu et al., 2024; Li et al., 2024; Zhuge et al., 2024). In this paper, we attempt to leverage large vision-language models (LVLMs) to model the practical court decisions on substantial similarity. However, directly applying large models for infringement identification may face unreliable outputs due to their limited comprehension or potential misinterpretation (Xu et al., 2025). To address this, we propose CopyJudge, an automated abstraction-filtration-comparison framework with multi-LVLM debate to reliably follow the court decision process on identifying substantial similarity.
Specifically, referring to software abstraction test (Abramson, 2002), we decompose the image into different layers or elements, such as composition, color, and theme, to distinguish between the basic concepts of the image and its specific expressions. Then, we filter out parts that are not copyright-protected, such as public and functional expressions. Finally, we compare the filtered portions to assess whether there is substantial similarity. To enhance the reliability of the judgment, we employ a multi-agent debate (Du et al., 2023; Chan et al., 2023) method where multiple LVLMs discuss and score the similarity. Each LVLM can make judgments based on the scores and reasons provided by other LVLMs. Ultimately, another LVLM-based meta-judge gives the final score and rationale based on the consensus of the debate. To enhance the consistency with human preferences, we inject human priors into each LVLM via few-shot demonstrations (Agarwal et al., 2024).
Given the judging results, we further explore how to mitigate infringement issues in text-to-image diffusion models. Utilizing our CopyJudge, we propose an easy but effective black-box infringement mitigation strategy that leverages a defense LVLM to iteratively optimize the input infringing prompts. This process avoids generating sensitive infringing expressions by querying supporting infringement rationales, while preserving the integrity of the original content. Moreover, if the input latent of the diffusion model is controllable, we could further enhance the mitigation approach by exploring specific non-infringing noise vectors within the latent space in a reinforcement manner, with the reward being reducing the predicted infringement score. This helps avoid infringement while maintaining the desired output characteristics, even without changing the original prompts.
@@ 2. Related Work
Copyright infringement in text-to-image models. Recent research (Carlini et al., 2023; Somepalli et al., 2023a, b; Gu et al., 2023; Wang et al., 2024d; Wen et al., 2024; Chiba-Okabe, 2024) highlights the potential for copyright infringement in text-to-image models. These models are trained on vast datasets that often include copyrighted material, which could inadvertently be memorized by the model during training (Vyas et al., 2023; Ren et al., 2024). Additionally, several studies have pointed out that synthetic images generated by these models might violate IP rights due to the inclusion of elements or styles that resemble copyrighted works (Poland, 2023; Wang et al., 2024b). Specifically, models like stable diffusion (Rombach et al., 2022) may generate images that bear close resemblances to copyrighted artworks, thus raising concerns about IP infringement (Shi et al., 2024b).
@@
Image infringement detection and mitigation. The current mainstream infringing image detection methods primarily measure the distance or invariance in pixel or embedding space (Carlini et al., 2023; Somepalli et al., 2023a; Shi et al., 2024b; Wang et al., 2021, 2024c). For example, Carlini et al. uses the $L_{2}$ norm to retrieve memorized images. Somepalli et al. use SSCD (Pizzi et al., 2022), which learns the invariance of image transformations to identify memorized prompts by comparing the generated images with the original training ones. Zhang et al. compare image similarity using the LPIPS distance, which aligns with human perception but has limitations in capturing certain nuances. Wang et al. transform the replication level of each image replica pair into a probability density function. Studies (Wen et al., 2024; Wang et al., 2024a) have shown that these methods have lower generalization capabilities and lack interpretability because they do not fully align with human judgment standards. For copyright infringement mitigation, the current approaches mainly involve machine unlearning to remove the model’s memory of copyright information (Bourtoule et al., 2021; Nguyen et al., 2022; Kumari et al., 2023; Zhang et al., 2024) or deleting duplicated samples from the training data (Webster et al., 2023; Somepalli et al., 2023b). These methods often require additional model training. On the other hand, Wen et al. have revealed the overfitting of memorized samples to specific input texts and attempt to modify prompts to mitigate the generation of memorized data. Similarly, Wang et al. leverage LVLMs to detect copyright information and use this information as negative prompts to suppress the generation of infringing images.
3. LVLM for Infringement Identification
@@
Figure 1. CopyJudge: an automated abstraction-filtering-comparison framework for image copyright infringement identification. This LVLM-based framework automatically decides whether a generated image infringes copyright by combining copyright expression extraction with infringement determination through multi-LVLM debate.
3.1. Problem Formulation
Use case. Our goal is to determine whether an image infringes the copyright of a known copyrighted image, where a copyright holder identifies a potentially infringing image and seeks to evaluate its similarity to their copyrighted work. In such scenarios, the copyright owner inherently possesses prior knowledge of the source material, making the requirement to specify it legally and practically justified. This aligns with real-world copyright enforcement workflows, where disputes often arise after a rights holder detects a suspected infringement and initiates a formal assessment.
@@ 3.1. Problem Formulation
Based on U.S. law (rot, 1970) and similar laws in other countries, given an image $x$ created with access to the copyrighted image $x_{cr}$, if $x$ and $x_{cr}$ exhibit substantial similarity, then $x$ is deemed to infringe the copyright of $x_{cr}$. Motivated by this, we aim to establish a substantial similarity identification model $f$, which takes $x$ and $x_{cr}$ as inputs and outputs a similarity score $s$. When $s$ exceeds a threshold $\gamma$, we determine that $x$ infringes on $x_{cr}$. This can be defined as:
(1) $\text{IsInfringement}(x)=\mathbb{I}(f(x,x_{{cr}})>\gamma),$
@@
3.2. Abstraction-Filtering-Comparison Framework
@@
[…]
(4) $\pi_{i}(x,x_{{cr}},z^{c},z^{c}_{cr},p_{i},D_{h},S_{h})\to(s_{i},c_{i},r_{i}).$
(5) $\left|s_{i}-s_{j}\right|\leq\alpha\quad\forall i,j\in\{1,2,\dots,N\}.$
Figure 2. Copyright infringement mitigation framework with controlling input prompts and latent noise. For prompt control, the defense agent iteratively optimizes prompts to mitigate copyright risks in generated images given the CopyJudge’s feedback. For latent control, it integrates a RL-based defense agent with reward-guided latent sampling to reduce the predicted infringement scores.
(6) $\pi_{f}(x,x_{{cr}},z^{c},z^{c}_{cr},p_{f},D_{h},S_{h},S_{m},C_{m},R_{m})\to(s_{f},c_{f},r_{f}),$
where $S_{m}$, $C_{m}$, and $R_{m}$ represent the set of scores, confidence levels, and rationales from $N$ LVLMs after reaching consensus in the $m$-th ($m\leq M$) debate. By combining the strengths of individual agents and iterative debating, the approach could achieve a reliable assessment of visual similarity.
Furthermore, we can determine whether the generated image constitutes infringement based on whether the final similarity score exceeds a specific threshold $\gamma$:
(7) $\text{IsInfringement}(x)=\mathbb{I}(s_{f})>\gamma.$
The whole two-stage process ensures a comprehensive and reliable evaluation by integrating multiple perspectives and rigorous analysis. Algorithm 1 shows the complete process. All instruction prompts can be found in the Appendix A.1.
4. LVLM for Infringement Mitigation
4.1. Mitigation via LVLM-based Prompt Control
(8) $\pi_{p}(x^{t},x_{{cr}},p^{t},p^{c},s^{t}_{f},c^{t}_{f},r^{t}_{f})\to p^{t+1}.$
Here, $p^{c}$ requires the modifier to alter the infringing expression while preserving the original expression as much as possible to avoid generating meaningless images. This mitigation strategy does not require any knowledge of text-to-image models, making it suitable for general black-box scenarios. The algorithm and instruction prompts can be seen in Algorithm 2 and Appendix B.1.